### v1

# Learning Attentive Neural Processes for Planning with Pushing Actions

Abstract
Our goal is to enable robots to plan sequences of tabletop actions to push a block with unknown physical properties to a desired goal pose on the table. We approach this problem by learning the constituent models of a Partially-Observable Markov Decision Process (POMDP), where the robot can observe the outcome of a push, but the physical properties of the block that govern the dynamics remain unknown. The pushing problem is a difficult POMDP to solve due to the challenge of state estimation. The physical properties have a nonlinear relationship with the outcomes, requiring computationally expensive methods, such as particle filters, to represent beliefs. Leveraging the Attentive Neural Process architecture, we propose to replace the particle filter with a neural network that learns the inference computation over the physical properties given a history of actions. This Neural Process is integrated into planning as the Neural Process Tree with Double Progressive Widening (NPT-DPW). Simulation results indicate that NPT-DPW generates more effective plans faster than traditional particle filter methods, even in complex pushing scenarios.
I Introduction
In scenarios where robots operate in unstructured, dynamic environments, planning under uncertainty is paramount. Some relevant properties for manipulation, such as object geometry, can be easily observed through sensors. Many other properties, such as inertial properties and friction, require interaction with the object [1]. One way robots can learn these inertial properties is to iteratively interact with objects and observe their behavior.
We focus on an instance of this problem where a robot must compose a sequence of pushes to guide a block from a starting configuration to a goal configuration, without prior knowledge of the center of mass. Instead, it must infer these properties while simultaneously learning a predictive model of the outcomes of pushing actions.
We model our planning challenge as a Partially Observable Markov Decision Process (POMDP), where unobservable physical properties remain constant over time [2]. Many POMDP solutions rely on Bayesian state estimation techniques (e.g., Kalman or particle filters) to update their beliefs as new observations become available. However, the linear assumptions required for Kalman filters do not hold in this scenario, since varying the center of mass can drastically alter rotational outcomes. More general methods, such as particle filters, must be used, despite their high computational cost [3].
To leverage these particle-based belief representations to solve our planning problem, we employ Particle Filter Trees with Double Progressive Widening (PFT-DPW), a Monte Carlo tree search algorithm that performs a randomized search to identify the plan yielding the highest expected reward. In POMDP planning, it is crucial to balance exploration and exploitation [4], because actions that best uncover hidden parameters may not directly move the block toward its goal. PFT-DPW and similar Monte Carlo tree search methods manage this trade-off effectively by systematically exploring uncertain aspects of the system while exploiting known information to achieve high rewards [5].
However, PFT-DPW also has significant limitations due to the use of a particle filter. First, updating the belief scales in cost linearly with the number of particles, which becomes prohibitively expensive when using large neural networks for each update. Since belief updates occur repeatedly as a subroutine within the planning algorithm, these costs accumulate quickly. Second, PFT-DPW is limited in its ability to handle more complex problems: it remains susceptible to the curse of dimensionality, where the required number of particles grows exponentially with the dimension of the underlying distribution, making it difficult to scale to higher-dimensional tasks.
To overcome these problems, we propose a model called the Pushing Neural Process (PNP), which jointly trains an inference network that estimates physical parameters, and a prediction network that learns to output a distribution of push outcomes based on physical properties. The inference network and prediction network are set up in an encoder-decoder architecture, where a representation of the physical properties is learned in the intermediate latent space.
We use our models to derive a new planner, which we call NPT-DPW (Neural Process Tree with Double Progressive Widening). NPT-DPW replaces the slower particle filter used to maintain belief in PFT-DPW with a neural network that consumes the history of a given search and outputs a unified belief distribution. Our results show that, on average, NPT-DPW reaches up to twice as close to the goal location as PFT-DPW and avoids catastrophic failures (such as pushing into prohibited zones) far more often.
Figure 1: Above is a diagram describing the Push Neural Process architecture and its use within NPT-DPW. On the right of the image, the search tree over actions and outcomes is pictured, where at each step an action is selected. We then pass in the history of outcomes ($h$) that we have already observed into the encoder ($q_{\phi}$) of the model to obtain the distribution over the latent parameters, which we sample to get ($z$). We then pass $z$ to the decoder with our new action $a$ to get a distribution over outcomes. We sample from that distribution as an outcome ($o$) for the action $a$. We then add that outcome to the potential outcomes in the search tree.
II Problem Setting
For the objects we aim to push, we categorize their properties into two distinct sets: observable properties, denoted by $x\in\mathcal{X}\subset\mathbb{R}^{n}$, and latent physical properties, denoted by $z\in\mathcal{Z}\subset\mathbb{R}^{m}$, where $m$ is decided empirically based on what results in the strongest model. The observable properties $x$ include features such as the block’s resting orientation, while $z$ captures physical properties such as center of mass. These latent properties are unknown to the robot a priori but play a critical role in determining the outcome of a push. We define a push in terms of an action $a\in\mathcal{A}$, characterized by an approach angle and a push velocity, where the angle is in the range $[0,2\pi]$. Upon completing a push, the robot observes the resulting pose of the block $o\in SE(2)$.
Within our Partially Observable Markov Decision Process (POMDP) framework, the transition function $T$ models the object’s dynamics. It maps the current state—which encompasses both the object’s observable pose and its latent physical properties—and an action to a probability distribution over future object poses. Importantly, the object’s pose remains observable to the robot at all times, while the latent properties remain constant throughout the interaction.
Consequently, our observation function $O$ simply reflects the observable portion of the state provided by the transition function $T$. The reward function $R$ penalizes the robot proportionally to its distance from the target goal, offering a substantial reward upon successfully reaching the goal (see Section IV-A, Equation 1 for explicit details). A discount factor $\gamma$ further encourages shorter plans aimed at efficiently reaching the goal.
Collectively, we formulate our pushing task as a POMDP defined by the tuple
$(\mathcal{S},\mathcal{A},\Omega,T,O,R,\gamma)$
where $\mathcal{S}=\mathcal{X}\times\mathcal{Z}$ represents all possible states and $\omega\in SE(2)$ represents all possible observations.
III Methods
III-A Pushing Neural Process
To solve the POMDP, we require probabilistic models that describe transitions in the partially observable state of the system and the observations made by the robot. We use an Attentive Neural Process, which we call a Pushing Neural Process, to jointly model push state estimation and prediction [6]. This model consists of two primary components: an encoder that infers latent representations from historical data, and a decoder that predicts future push outcomes given these representations.
We denote a sequence of pushes and their outcomes as the dataset:
$D_{t}=\{(a_{1},o_{1}),\ldots,(a_{t},o_{t})\}.$
The encoder consumes $D_{t}$ (of arbitrary length) along with the observable properties $x$ (object geometry) and infers a probability distribution over the latent parameters $z$. To this end, we train a neural network, parameterized by $\phi$, to produce an approximate posterior $q_{\phi}(z|x,D_{t})$ that closely matches the true posterior $p(z|x,D_{t})$.
Given the latent representation from the encoder, the decoder—parameterized by $\theta$—then predicts push outcomes. Mathematically, prediction of a new outcome $o$ given action $a$, object properties $x$, and history $D_{t}$ can be expressed using the law of total probability:
[…]
Algorithm 1 NPT-DPW
procedure SimulateAction($h,a$)
     $b\leftarrow$PNPEncoder(h) $\triangleright$ PNP inference step.
     $p\leftarrow$ PNPDecoder(b, a) $\triangleright$ PNP prediction step.
     $o\leftarrow p.sample()$
     $r\leftarrow\text{Reward}(o.x,o.y)$triangleright As defined in Equation 1
     return $o,r$
end procedure
procedure Simulate($h,d$)
     if $d=0$ then
         return 0
     end if
     $a\leftarrow$ ActionProgWiden(h)
     if $|C(ha)|\leq N(ha)^{\alpha_{0}}$ then
         $o,r\leftarrow$ SimulateAction(h,a)
         $C(ha)\leftarrow C(ha)\cup\{o\}$
         $total\leftarrow r+\gamma\cdot$Rollout($hao,d-1$)
     else
         $o,r\leftarrow$ sample from $C(ha)$
         $total\leftarrow r+\gamma\cdot$Simulate($hao,d-1$)
     end if
     return $total$
end procedure
Since a single sample from $p(z|x,D_{t})$ is an unbiased estimator of the distribution, we sample a single $z\sim q_{\phi}(z|x,D_{t})$. The decoder consumes $z$, the observable object properties, and the new push to predict a distribution over the observation (block displacement) $p_{\theta}(o|a,x,z)$.
During training, we maintain a collection of pushes for each object in a set of objects with varying physical properties. We provide the partial dataset $D_{t}$ as input to the inference network (encoder), and apply the prediction network (decoder) to the collection of all pushes $D_{T}$ for the given object:
$D_{T}=D_{t}\cup\{(a_{t+1},o_{t+1}),\ldots,(a_{T},o_{T})\}.$
We optimize the Evidence Lower Bound (ELBO)
$\mathbb{E}\Bigl{[}\log p_{\theta}(o|x,a,z)\;-\;D_{KL}\bigl{(}q_{\phi}(z|D_{T})\;\|\;q_{\phi}(z|D_{t})\bigr{)}\Bigr{]},$
which promotes accurate push outcome predictions while reducing the discrepancy between the full and partial posterior distributions. The cross-entropy term increases the likelihood of accurately predicting outcomes over time, and the KL term mitigates overfitting.
III-B NPT-DPW
Our planner is an extension of PFT-DPW introduced by Sunberg and Kochenderfer [5]. In PFT-DPW, a particle filter is used to represent the set of hypotheses over the partially observable state of the problem, and the belief is updated as new push outcomes are observed. The PF is used as a belief state in Monte Carlo tree search, which branches at new actions and observations. Since both the action and observation spaces are continuous, double progressive widening is used to limit the number of actions and observations sampled at each tree node to ensure reliable reward estimates.
Instead of using a particle filter, our planner, NPT-DPW, uses the Pushing Neural Process to infer the underlying physical parameters and predict outcomes of future actions. To expand the tree, we pass the history of actions and outcomes through the encoder to infer the physical parameters. Actions are sampled as they normally would be in PFT-DPW. Observations are sampled from the distribution computed by the decoder from the inferred parameters and the push under consideration. Thus, every step of the tree search will have runtime complexity of $O(h^{2})$, where $h$ is the size of the history (due to the quadratic complexity of self-attention), rather than $O(n)$, where $n$ is the number of particles used in PFT-DPW. We hypothesize a significant improvement in plan quality within a fixed computational budget, since $h\ll n$ in practice.
The NPT-DPW algorithm is described in Algorithm 1; any functions not described in this block are the same as those specified for PFT-DPW in Sunberg and Kochenderfer [5]. Furthermore, a diagram of the neural process and its integration into the planner can be seen in Figure 1.
Figure 2: The pushing surfaces used in our experiments. The rectangle denotes the starting position, and the star the goal.
IV Experiments
Figure 3: Results plotting average distance from the goal as a function of planning time allocated per action. It demonstrates the effectiveness of the NPT-DPW versus PFT-DPW (with 10/30/100 particles). We conducted trials with 0.5,1,2,3,5, and 10 seconds of time to plan per step.
In our experiments, we seek to evaluate the performance of NPT-DPW relative to the baseline method PFT-DPW, given a fixed computational runtime budget. The problem scenarios are designed to evaluate whether an MCTS-DPW planning scheme can leverage the efficient inference updates of the Neural Process for a deeper tree search, yielding better plans.
Figure 4: View of the robot interacting with the block in simulation.
IV-A Scenario
The experiments will be run using three tables with geometry designed to vary the hardness of the pushing problem (Figure 2). Each of these tables defines a different planning problem. The first is a blank-surface table designed to check whether the planner can produce plans that drive the block to the goal. In the second and third scenarios, the movement of the block must be constrained along certain paths to prevent it from falling off the pushing surface. The second one is slightly simpler since it encourages pushing in a straight line, while the third will require the robot to understand how to rotate the block. Solving these instances requires the planner to accurately infer the object’s physical properties. An intelligent planner must sequence information-gathering actions on a wider area before pushing the block along the narrow sections of these tables.
After each push, the robot obtains a reward for the block’s location $(x,y)$ relative to the goal $(g_{x},g_{y})$. We choose the reward function
$\text{Reward}(x,y)=$ $-10\sqrt{(x-g_{x})^{2}+(y-g_{y})^{2}}$
$-100\cdot\texttt{fallen}(x,y)$
$+10000\cdot\texttt{inRange}(x,y)$ (1)
fallen is a function that returns 1 if the block is off the table, and 0 otherwise. inRange is a function that is $1$ if the block is within 0.2 meters of the goal, and 00 otherwise. We set the discount factor $\gamma=0.6$ to incentivize the planner to prioritize not falling off the table in the next action. An example of a plan created by PFT-DPW is pictured in Figure 5.
We assess performance by fixing the computation time $t$ allowed at each step of the plan. Then for each $t$ in $[0.5,1,2,3,5,10]$ seconds, we ran 15 trials on each map for each planner, where we recorded the distance of the object from the goal after 30 steps. If the object fell off the table, or reached the goal early, we stopped then. We sampled a new center of mass for the block in each trial.
We also varied the number of particles in PFT-DPW, initializing it with $n=10,30$, or $100$, to evaluate NPT-DPW against a range of particle filter–based planners and assess the impact of particle count on performance.
These simulations will be run in the PyBullet simulator (Figure 4), with the robot being programmed to push in a direction while keeping its height and direction consistent throughout the motion. To sample actions for progressive widening, we uniformly sample an angle from $[0,2\pi]$ in world coordinates and push at the velocity of $0.10$ m/sec. Each of the blocks is a rectangular prism with a fixed geometry of 2.5 cm x 2.5 cm x 1.5 cm. The mass of the block is set constant at 0.2 kg. The center of mass of the block is sampled uniformly within the convex hull of the object’s geometry.
Our neural network architecture is based on Attentive Neural Processes [6]. Each push is first encoded by an MLP and processed through multiple layers of self-attention. After that, the outputs of the self-attention layers are aggregated into outputs consisting of a mean and covariance. These represent the mean and diagonal covariance of a multivariate Gaussian over $z$, representing the distribution over latent physical properties. Note that for our problem, the only latent physical properties are the $x$ and $y$ coordinate of the center of mass, but we still set $z\in\mathbb{R}^{5}$ since it empirically led to improved performance. The decoder is structured as multiple linear layers that consume a sample from the latent space and outputs a distribution over possible outcome poses of the block after the push is finished.
Figure 5: This is an example of a plan created by PFT-DPW and run by the simulation on the ”ring” map. Important features that show signs of intelligent planning are the pushes in the beginning in the corner, signifying learning more about the block before trying to move towards the goal. Note the figure has been re-scaled slightly from the real experiments to be more visually clear.
IV-B Results
Our results are shown in Figure 3. When the planner is allocated a second or more to plan, NPT-DPW significantly outperforms PFT-DPW. On the ”beam” map, NPT-DPW performs more than twice as well on average. However, when only half a second per action is allocated, the performance of both models is similarly poor. This is likely because the planners may fail to sample an action angle that moves the block toward the goal.
A plausible explanation for NPT-DPW’s better performance quicker belief updates that allowed for more traversals of the tree. For instance, NPT-DPW always ended up searching at least twice as many actions as PFT-DPW. Additionally, we hypothesize that the self-attention heads in the encoder of the Attentive Neural Process help the model learn structural information about the block. This is because self-attention allows the model to infer relationships between previous push outcomes. In contrast, the particle filter maintains a belief that is updated sequentially with each new observation, which still allows for learning interdependence, but doesn’t directly compare pushes like the neural process does.
We believe the lack of performance differences between various particle counts in PFT-DPW can be attributed to the nature of the problem being solved. In our simulation, small differences in the center of mass are difficult to distinguish, since the push outcome is primarily affected by the relative location of the center of mass to the point of contact. As a result, using only a few particles may yield most of the benefits that additional particles provide, while also allowing the planner to explore more push actions due to cheaper belief updates. This trade-off may explain the trends observed in Figure 3. We expect that with planning times longer than those tested, particle filter–based planners with more particles may eventually surpass smaller models in performance.
Ultimately, our results suggest the viability of using an Attentive Neural Process as an alternative to traditional particle filter–based approaches for planning in POMDPs. In particular, when given less than 10 seconds to plan, NPT-DPW consistently and significantly outperforms PFT-DPW.
V Related Work
Our pushing POMDP is an instance of a Hidden-Parameter POMDP [2], where the unobserved state variables remain static but influence the transition dynamics of the planning problem. Several previous works have introduced Bayesian models for rapid estimation and adaptation of the unobserved state in this class of problems. The difference between these methods largely depends on the representation of uncertainty. Killian et al. [7] use Bayesian neural nets; others use Gaussian processes [2, 8]. Ensembles of neural networks are a valid choice as well [9].
Other work has focused on learning object dynamics with unobservable object properties outside the POMDP framework. Many authors develop ways to estimate unobserved properties when the dynamics are known [10, 11, 12, 13]. Other works, like our method, do not assume the dynamics are known [14, 15, 16, 17]. However, these works assume they have access to a dataset of interactions from a new instance of the system to estimate the unobserved variables.
Estimating unobserved properties through interaction is a common strategy in many manipulation tasks [1]. These strategies are applied to adapt grasping classifiers [18, 19, 20, 21, 22], disentangle object geometry without visual input [23, 24], and obtain object articulation constraints [25, 26]. Bauza and Rodriguez [27] develop an alternate method to compute uncertainty of push outcomes based on uncertainty of physical parameters. Many of these methods tackle information gathering by developing problem-specific solutions [21, 22], or relying on bandit-style exploration approaches [19, 20] or information-gain heuristics [18, 24] for short, single-step estimation tasks. Our work leverages POMDP solvers to tackle the information-gathering problem when multiple actions must be composed together.
VI Conclusion
We present: the Pushing Neural Process, an Attentive Neural Process architecture to learn push outcome prediction of a novel object instance, and NPT-DPW, a new planner that integrates the PNP to produce good plans under reasonable computational runtime budgets.
Our initial experiments demonstrate promising capabilities; however, we acknowledge certain limitations that we aim to address in subsequent research. For instance, the PNP exhibited underfitting, which we attribute to the complexity and inherent instability of the simulation environment used to generate the training data. To address this, we hope to apply the NPT-DPW to a broader spectrum of tasks, including scenarios with simpler dynamics and transition models to yield clearer insights. Secondly, we observed occasional erratic behavior in generated plans, indicating opportunities for further model improvement. Addressing these issues may improve planning performance in both PFT-DPW and NPT-DPW. Finally, we hope to deploy the PNP on a physical robot platform in order to validate our approach more comprehensively and demonstrate applicability.
ACKNOWLEDGMENTS
We thank Michael Noseworthy and Ben Zandonati for their helpful feedback and support.